$ 18:50: Welcome to LBT, Dr. Sand! High winds, but it should be okay.

$ 19:10: Twilight flats

Filter scale counts
U 1.0 30000, 25000, *
V 20.0 27000, **, 8000, 6500, 5500

*: CCD communication error required LBC kill/start. Skipped remaining Uspec flats
**: Shutter locked up, fixed with testpower script on CMU

Judging from the stars in the flats, the seeing is probably lousy. Guessing 1.5 arcsec.

$ 19:45: Focusing at mpobj3_2 position
Took a while to find focus, required manual intervention

$ 20:08: mpobj3_2_g -- seeing 1.2--1.4 arcsec.
We did mpobj3_1 on 20070412, so these observations will complete mpobj3.

$ 20:33: mpobj3_2_r -- Shutter locked again. This time testpower script did not fix it. We tried a couple of times.
Now going to reboot the CMU.

$ 21:00: It's been a while, so focusing again. All right, the shutter is operational again.

$ 21:04: mpobj3_2_r -- First image has artifacts that look to me like the shutter was stuck open during readout. We're letting the second exposure continue. Seeing now about 1.7

$ 21:20: mpobj3_2_r -- Tried fixing shutter again with the testpower script. Restarting the OB.
The shutter is still stuck open. We are consulting John Hill.

$ 22:15: We went up to the camera and Olga tried toggling some switches. We'll try a test exposure to see what happened.

$ 22:43: After another CMU reboot, the shutter finally seems to be functioning! Trying a focus run at mpobj3 Darn. It's a bias. Olga tried the testpower script again, and it seemed to work, we were able to focus (three iterations).

$ 22:52: mpobj3_2_r (once again into the breach) -- It worked! Seeing 1.2-1.5 arcsec

$ 23:17: Focusing at mpobj2_1_g
First exposure: guider failed. Seeing was 1.8--2.0 arcsec
We started the second one, and guiding is working on this one. Seeing is still 1.8 arcsec.

There are no targets in the queue for 1.8 arcsec. David is looking at some proposals, to see if any of them can use data with 1.8 arcsec seeing. Meanwhile, we're going to take a dithered image of the coma cluster, and if the seeing improves, we'll immediately go back to the queue.

$ 00:19: Coma cluster imaging. Seeing in first exposure is 1.65 arcsec...since this seeing triggers proposals with 1.5 arcsec, we're going to go ahead and try some of those.

$ 00:24: Focus at position of mpobj2_1_g

$ 00:31: mpobj2_1_g -- Well, the seeing still sucks. We're just going to stick with these OBs. If nothing else, they can be used for dark-sky flats.....

Seeing 1.9 arcsec

Exp No. 2 is 1.9 arcsec

Exp No. 3 -- guiding failed! Well, this is because the shutter did not open.

Tried running the testpower script. No luck.

Tried restarting LBC. Ok, it worked. However, we seem to have lost exposure 3.

Doing Exp. 4 -- Shutter didn't close...actually, it now appears to be stuck partway open, because the images have an odd illumination pattern (combination of vignetting and scattered light?)

We are going to call Italy for advice. They will do tests on the camera tomorrow.

We cycled power on LBC, then tried a focus image, but it showed the shutter stuck open.

Now trying to reboot the CMU.

Fernando says the shutter is mechanically jammed. It's going to require a fix by the day crew.

-- JasonHarris - 16 Apr 2007
Topic revision: r3 - 16 Apr 2007, JasonHarris
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback